OpenStudy (anonymous):

lim x->16 (4-sqrt(x))/(s-16)

OpenStudy (psymon):

is it supposed to be s on bottom?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[\frac{ 4-\sqrt{s} }{ s-16 }\]

OpenStudy (psymon):

Oh, so its supposed tobe as "s" goes to 16, not x.

OpenStudy (anonymous):

oh yeah, sorry

OpenStudy (psymon):

No guarantees itll work, but when you see a square root +/- something, its best to try multiplying top and bottomby the conjugate first.

OpenStudy (anonymous):

Yeah, i did that. I just don't know how to factor out (16-s)/(s-16)(4+sqrt(s))

OpenStudy (inkyvoyd):

for (16-s)/[(s-16)(4+sqrt(s))] simply distribute the denominator as you would any algebraic expression, and then see what further simplification you can do.

OpenStudy (inkyvoyd):

an easier way would be to factor out the bottom as well

OpenStudy (psymon):

Hmm...maybe not/

OpenStudy (psymon):

\[\frac{ -(s-16) }{ (s-16)(4+\sqrt{s}) }\]

OpenStudy (anonymous):

that'd still be dividing by 0

OpenStudy (psymon):

Because (16-s) = (-s+16) = -(s-16)

OpenStudy (anonymous):

so factor out the negative from top then, not denominator

OpenStudy (psymon):

Doesnt matter honestly.

OpenStudy (inkyvoyd):

\(\huge \frac{(4-\sqrt s)}{(s-16)}=\frac{4-\sqrt s}{(\sqrt s-4)(\sqrt s+4)}=-\frac{\sqrt s-4}{(\sqrt s-4)(\sqrt s+4)}\)

OpenStudy (psymon):

\[\frac{ (16-s) }{-(16-s)(4+\sqrt{s}) }\]

OpenStudy (anonymous):

i'm getting -1/8

OpenStudy (psymon):

Yep :3

OpenStudy (psymon):

And what inky did is fine, too, he just did a difference of squares to factor it.

OpenStudy (anonymous):

so pretty much with limits; anything you can think of to make it not 0/0 or x/0

OpenStudy (psymon):

Basically. There are a lot of little tricks you can do, but if I ever see sqrt +/- something then conjugate is first thought.

OpenStudy (inkyvoyd):

The problem of having to rationalize the denominator is classic for limits. It pops up almost any class to give the problem of needing to multiply by the conjugate - little algebra fun that everyone loves

OpenStudy (psymon):

Sometimes you may have to rationalize the denominator, sometimes the numerator, sometimes even both at the same time.

OpenStudy (anonymous):

Yeah, I just didn't think to factor out the negative.

OpenStudy (inkyvoyd):

you will learn, in the future, ways to resolve indeterminate expressions for limits - for now though, stick to simplification and trying get rid of determinant forms

OpenStudy (psymon):

Yeah, calc 2 gives ya a trick to deal with anything that ends up as 0/0 or infinity/infinity.

OpenStudy (anonymous):

I have another one if you guys are interested.

OpenStudy (psymon):

Go for it.

OpenStudy (inkyvoyd):

Post as new question :)

OpenStudy (anonymous):

lim x ->9 sqrt(x)/(x-9)^4

OpenStudy (anonymous):

oh okay i will